OpenStudy (anonymous):

What is the slope of the line that passes through the given points? (6, 2) and (7, 4)

hartnn (hartnn):

slope of the line that passes through (x1,y1)and(x2,y2) is m = \(\frac{y2-y1}{x2-x1}\) does this help ?

OpenStudy (anonymous):

Take 6 as x1, 2 as y1 and 7 as x2, 4 as y2 Then apply the formula, m/slope = y2-y1/(x2-x1) and substitute the formula...what do you get ? :)

OpenStudy (anonymous):

The options are A. -2 B. 1/2 C. -(1/2) D. 2

OpenStudy (anonymous):

4-2/(7-6) = ?

OpenStudy (anonymous):

D?

hartnn (hartnn):

yup :)

OpenStudy (anonymous):

Exactly! You're awesome :D

OpenStudy (anonymous):

Thank you both.

hartnn (hartnn):

welcome :)

OpenStudy (anonymous):

Our pleasure xD